OpenStudy (anonymous):

state whether the given measurements determine zero one or two triangles. C = 30° a = 28 c = 14

OpenStudy (anonymous):

USE SIN FORMULA to find side a.Namely sinA/a=sinB/b=sinC/c

OpenStudy (anonymous):

sinA=(a/c)sinC. sinA=1. A=90 Degree

OpenStudy (anonymous):

use Pythagoras theorem to find sides